# {textblock*}

### v1

@@ 2 Preliminary Reflections on the Horizon Criterion
As a result, the most interesting situations to study are ETW branes bounding spacetimes without Minkowski or AdS asymptotics. Concretely, one may think of spherical ETW branes bounding ball-shaped universes created from nothing via a ‘bubble of something’ [79, 80, 81, 63] process. However, we will soon forget about the origin of such universes and focus on their asymptotically late phase and the conditions for brane-brane horizons to exist.
@@
Before analyzing in detail the ‘brane-brane horizon criterion’ as stated in the Introduction, we want to mention a possible alternative conjecture: We could try to forbid the eternal existence of spatially compact universes. Indeed, when applied to spheres, this would have implications very similar to the horizon criterion. This can be understood as follows: Consider the Hartle-Hawking or Linde-Vilenkin creation-from-nothing process [82, 83, 84, 85]. The resulting spatial spheres are driven to collapse by their curvature and to expansion by their vacuum energy. Demanding that their lifetime is finite will then place a abound on how flat asymptotically positive potentials may be, similarly to the horizon criterion. However, there is a catch: As noted long ago [86, 87, 88], also spatially flat (e.g. toroidal) or negatively curved compact universes can form spontaneously. As there is no curvature effect favoring collapse, they can live eternally, even at vanishing vacuum energy. Hence the eternal existence of compact universes can not be the fundamental issue responsible for the problems with dS.
@@
must hold at late times [36]. For FLRW universes, this is equivalent to $\ddot{a}(\tau)\leq 0$.
@@
Requiring (5) or $A_{\mu}X^{\mu}\leq 0$ to hold for $\tau>\tau_{0}$ provides a local condition on the spacetime geometry that forbids eternal accelerated expansion. However, there exist constructions in string theory where $\ddot{a}(\tau)$ approaches zero from above for $\tau\to\infty$ [49, 50, 4031], such that this condition is too strong. At best, one could demand that the limiting value of $\ddot{a}$ at $\tau\to\infty$ is non-positive. Such a statement is then somewhat unsatisfying as there is no quantification of how quickly $R_{\mu\nu}v^{\mu}v^{\nu}$ or $S(A^{\mu})$ must approach zero. To exclude cosmologies where $\ddot{a}$ decays arbitrarily slowly, [30] proposes that the equation of state parameter $w$ must asymptote to its critical value or exceed it. But this is of course somewhat involved.
We see that appropriate conditions on $R_{\mu\nu}v^{\mu}v^{\nu}$ or $S(A^{\mu})$ probably have to involve a nontrivial specification of how fast this quantity must decay asymptotically. This appears ad-hoc and it thus remains unclear whether a convincing constraint involving locally defined quantities can be found. We conclude that the presence of horizons, as suggested in this context in [30, 31], remains in many ways the conceptually simpler and hence hopefully more fundamental property to consider.
@@ 3 ETW Brane with brane-localized Scalar
We are interested in solutions where a homogeneous and isotropic ETW brane asymptotically expands. This of course includes different sub-cases: The brane could be spatially flat or possess negative or positive curvature. Similarly, the adjacent bulk could also be curved and the curvature could even vary transversely to the wall without reducing the symmetry of the solution. Clearly, all such curvature effects will asymptotically decay and we expect that picking one of the many possibilities outlined above is sufficient to derive the general bound on $V(\phi)$. This expectation is based on the fact that, in the boundary-free situation, it is known that curvature effects are insufficient to induce horizons if the scalar potential is too steep to do so [31]. Nevertheless, this is of course not a proof and further scrutiny may be warranted. Based on these preliminary remarks, we first study spherical ETW branes bounding a solid cylinder of Minkowski space, with the cross section being a ball. Based on the intuition developed, we then proceed by studying a setting where the ETW brane is embedded with flat spatial sections.
@@ 3.1 Spherical ETW Brane
The equations of motion following from the action (9) are [^7] These equations can be obtained from the Israel junction conditions $M_{d}^{d-2}(K_{ab}-h_{ab}K)=S_{ab}$ [89], where $K_{ab}$ is the extrinsic curvature and $S_{ab}$ is the energy momentum tensor of the fields living on the boundary. The extrinsic curvature is defined as $K_{ab}=\nabla_{a}n_{b}$, where $n$ is the outward-pointing unit normal vector of the spacetime boundary.
      Note that our equations differ from much of the brane-cosmology literature (see e.g. [56, 90, 91, 92, 93]) as these papers work locally on the covering space of a $\mathbb{Z}_{2}$ orbifold. By contrast, we treat the brane locally as a boundary of a half-space. Agreement may be achieved through the identification $\rho_{\rm there}=2\rho_{\rm here}$.
$H^{2}=\frac{M_{d}^{-2(d-2)}}{(d-2)^{2}}\rho^{2}-\frac{1}{R^{2}}\,,\qquad$ $H\equiv\frac{\dot{R}}{R}\,\,\,,$ (13)
$M_{d}^{d-3}\left(\ddot{\phi}+(d-2)H\dot{\phi}\right)+\partial_{\phi}V(\phi)=0\,,\qquad$ $\rho\equiv\frac{1}{2}M_{d}^{d-3}\dot{\phi}^{2}+V(\phi)\,\,,$ (14)
and the ‘dot’ stands for $d/d\tau$. Furthermore, it follows from the derivation of (13) that the condition $\rho>0$ must hold. Equation (13) is similar to the Friedmann equation for a closed universe, but with a key difference [90, 91]: The energy density $\rho$ appears quadratically rather than linearly, implying a substantially different dynamics.
A necessary condition for the existence of brane-brane horizons in the $d$-dimensional spacetime is that the induced, $(d\!-\!1)$-dimensional geometry of the boundary has horizons. The condition for these to occur is well-known: It is given by (2) where the scale factor is identified with $R(\tau)$. Note that only the asymptotic form of $R(\tau)$ is of relevance. It follows that if
@@
In the presently discussed situation, the condition above is not sufficient: Indeed, two points $p,p^{\prime}$ on the boundary can be causally connected in the full $d$-dimensional spacetime but not in the $(d\!-\!1)$-dimensional geometry of the boundary. A sufficient condition for brane-brane horizons to exist is the following: An observer on the ETW brane should not be able to send a signal to the antipodal point of the spherical brane, not even through the bulk. This corresponds to demanding that
@@
[…]
@@ 3.2 ETW Branes with flat Slicing
We believe the spherical case discussed above to be the most intuitive and technically simplest. Nevertheless, as just explained, one could suspect that a flat ETW brane embedding leads to a stronger condition. To clarify this, we build on [94], but generalize the analysis to arbitrary dimensions. The metric ansatz is
$\begin{split}ds_{\rm g}^{2}&=e^{2A(t,z)}(-dt^{2}+dz^{2})+e^{2B(t,z)}ds_{d-2}^{2}=-e^{2A(u,v)}dudv+e^{2B(u,v)}ds_{d-2}^{2}\,,\end{split}$ (21)
where $u=t-z$, $v=t+z$ and $ds_{d-2}$ is the spatial metric of the ETW brane. The latter is taken to be localized at $z=0$ and the bulk is the region $z>0$. The metric (21) follows from considering an FLRW-type brane embedding and choosing conformal gauge in the directions transverse to $ds_{d-2}^{2}$. The induced metric on the worldvolume of the ETW brane can be written as
@@
with $\rho$ and $p$ being the energy density and pressure on the ETW brane induced by the field $\phi$. We now make an important assumption: We consider $A$ and $B$ to depend on $u=t-z$ only. One may understand this requirement as choosing the boundary conditions such that no incoming flux of energy hits the ETW brane. Furthermore, this assumption ensures that the bulk geometry is Riemann-flat. It follows that there remains a single nontrivial Einstein equation (25) to be solved. It can be integrated to yield [^9] The integration constant can be absorbed into a redefinition of $u$.
@@
We thus find, in accordance with [91, 95],
@@
the Friedmann equation of type (13). Because of the flat brane slicing the curvature term $R^{-2}$ is absent. Equation (32) and the scalar field equation (14) determine the evolution of the system. Our assumption that $A$ and $B$ depend only on $u=t-z$ is related to setting the projected 5d Weyl tensor $E_{ab}$ to zero, see e.g. [91, 95]. We note that the second Israel condition (29) is a consequence of (28) and the the scalar field equation (14).
@@
Importantly, this shows that for $\alpha>2$, the coordinate range of $t$ at the ETW brane is bounded by some constant $u_{0}$. In other words, $\tau\to\infty$ corresponds to $t\to u_{0}$. By contrast, for $21<\alpha<$ 2, $\tau\to\infty$ corresponds to $t\to\infty$. Furthermore, it is shown in App. A that the functions $e^{A(u)}$ and $e^{B(u)}$ behave as
$\begin{split}e^{A(u)}&=\alpha\left[\alpha(2-\alpha)(u-u_{0})\right]^{\frac{\alpha-1}{2-\alpha}}\,,\\ e^{B(u)}&=\left[\alpha(2-\alpha)(u-u_{0})\right]^{\frac{\alpha}{2-\alpha}}\,,\end{split}$ (34)
@@
[…]
@@
Let us now consider the case $21<\alpha<$ 2, for which $t$ is unbounded from above. In terms of the variable $u(t)=t-z(t)$, the integral (36) reads
[…]
@@
The corresponding coordinate transformation is derived in App. A.2 for the case $21<\alpha<$ 2. There, it is also shown that the ETW brane embedding is specified by the equation
@@ 3.3 Implications for the Potential
This is known in the case $d=5$, where the corresponding scaling solutions have been studied in the context of brane cosmologies [92] (see also [93, 96]).
@@
Equations (53)–(56) together with the definition (47) are the first consequences of the brane-brane horizon criterion.
@@
As a final comment, let us briefly think about explicit stringy examples for brane-localized scalars: One class of models are those of ‘Verlinde type’ [67] (mentioned in the Introduction), where a 6d compact space realizes the 4d ETW brane (concretely a $T^{6}/\mathbbm{Z}_{2}$) and a 5d compact space (concretely a fluxed $S^{5}$) realizes the bulk by dimensional reduction. The brane scalar would be the volume of the $T^{6}/\mathbbm{Z}_{2}$. Crucially, if this volume rolls to infinity, higher-dimensional brane-localized operators (in our case a 4d Einstein-Hilbert term) may come to dominate. This situation could arise fairly generally since infinite distances are almost always related to some type decompactification [97]. Thus, more work is required to make contact between our simple analysis and string models. Alternatively, if both bulk and brane decompactify or roll to weak coupling together, the above issue is avoided. Thus, we now turn to the case with a bulk scalar field which, as we just argued, may be the more interesting one from the string-theory perspective.
@@ 4 ETW brane coupled to massless Bulk Scalar
As already explained in the Introduction, one may think of $\phi$ as a bulk modulus. The brane generically has less supersymmetry and it is hence natural to expect that it provides a non-trivial potential depending on $\phi$. We do not include a brane-localized kinetic term for $\phi$: While such a term is generically present, its effects are sub-leading [98]. [^11] On the one hand, brane-parallel spatial gradients decay quickly as the brane expands. On the other hand, the kinetic energy $\dot{\phi}^{2}$ receives contributions from the brane and from the adjacent bulk region, with the latter dominating. This can be traced to a $1/M_{P}$ suppression arising on dimensional grounds [98].
@@ 4.1 The Equations of Motion
[…]
00 0 $=$ $\left(\left.-M_{d}^{d-2}n^{\mu}\partial_{\mu}\phi-\partial_{\phi}V\right)\right|_{\rm ETW}\,,$ (63)
where $\mathcal{K}_{d,\tau\tau}$ is the $\tau\tau$-component of the extrinsic curvature tensor. As before, $\tau$ labels the eigentime of the ETW brane such that $h_{\tau\tau}=-1$. The boundary condition (63) is the straightforward generalization of the familiar Neumann boundary condition to the case with a boundary-localized potential, see e.g. [99]. In the equations above, $n^{\mu}$ is the (outwards-pointing) unit normal vector of the spacetime boundary and $|_{\rm ETW}$ stands for evaluation at the ETW brane. As before, we focus on positive potentials that approach zero monotonically as $\phi\to\infty$.
We again build on [94] and consider a metric ansatz of the form (21). As before, we assume the functions $A$ and $B$ as well as the scalar field $\phi$ to depend on $u=t-z$ only. The single remaining nontrivial Einstein equation now takes the form

### v2

@@ 2.2 Potential counterexamples
As a result, the most interesting situations to study are ETW branes bounding spacetimes without Minkowski or AdS asymptotics. Concretely, one may think of spherical ETW branes bounding ball-shaped universes created from nothing via a ‘bubble of something’ Hawking:1998bn, Turok:1998he, Blanco-Pillado:2011fcm, Friedrich:2023tid process. However, we will soon forget about the origin of such universes and focus on their asymptotically late phase and the conditions for brane-brane horizons to exist.
2.3 Alternative formulations of a criterion against de-Sitter-like spacetimes
Before analyzing the Refined Horizon Criterion in detail, we use this subsection to critically discuss possible alternatives.
@@ 2.3 Alternative formulations of a criterion against de-Sitter-like spacetimes
We could try to forbid the eternal existence of spatially compact universes. Indeed, when applied to spheres, this would have implications very similar to the Horizon Criterion. This can be understood as follows: Consider the Hartle-Hawking or Linde-Vilenkin creation-from-nothing process Vilenkin:1982de, Hartle:1983ai, Linde:1983cm, Vilenkin:1983xq. The resulting spatial spheres are driven to collapse by their curvature and to expansion by their vacuum energy. Demanding that their lifetime is finite will then place a abound on how flat asymptotically positive potentials may be, similarly to the Horizon Criterion. However, there is a catch: As noted long ago Zeldovich:1984vk, Coule:1999wg, Linde:2004nz, also spatially flat (e.g. toroidal) or negatively curved compact universes can form spontaneously. As there is no curvature effect favoring collapse, they can live eternally, even at vanishing vacuum energy (see however Bedroya:2025ris). Hence the eternal existence of compact universes can not be the fundamental issue responsible for the problems with dS.
@@
must hold at late times Rudelius:2021azq. For FLRW universes, this is equivalent to $\ddot{a}(\tau)\leq 0$.
@@
Requiring (5) or $A_{\mu}X^{\mu}\leq 0$ to hold for $\tau>\tau_{0}$ provides a local condition on the spacetime geometry that forbids eternal accelerated expansion. However, there exist constructions in string theory where $\ddot{a}(\tau)$ approaches zero from above for $\tau\to\infty$ [Chen:2003dca, Andersson:2006du, Marconnet:2022fmxAndriot:2023wvg], such that this condition is too strong. At best, one could demand that the limiting value of $\ddot{a}$ at $\tau\to\infty$ is non-positive. Such a statement is then somewhat unsatisfying as there is no quantification of how quickly $R_{\mu\nu}v^{\mu}v^{\nu}$ or $S(A^{\mu})$ must approach zero. To exclude cosmologies where $\ddot{a}$ decays arbitrarily slowly, Hebecker:2023qke proposes that the equation of state parameter $w$ must asymptote to its critical value or exceed it. But this is of course somewhat involved.
We see that appropriate conditions on $R_{\mu\nu}v^{\mu}v^{\nu}$ or $S(A^{\mu})$ probably have to involve a nontrivial specification of how fast this quantity must decay asymptotically. This appears ad-hoc and it thus remains unclear whether a convincing constraint involving locally defined quantities can be found. We conclude that the presence of horizons, as suggested in this context in Hebecker:2023qke, Andriot:2023wvg, remains in many ways the conceptually simpler and hence hopefully more fundamental property to consider.
@@ 3 ETW Brane with brane-localized Scalar
We are interested in solutions where a homogeneous and isotropic ETW brane asymptotically expands. This of course includes different sub-cases: The brane could be spatially flat or possess negative or positive curvature. Similarly, the adjacent bulk could also be curved and the curvature could even vary transversely to the wall without reducing the symmetry of the solution. Clearly, all such curvature effects will asymptotically decay and we expect that picking one of the many possibilities outlined above is sufficient to derive the general bound on $V(\phi)$. This expectation is based on the fact that, in the boundary-free situation, it is known that curvature effects are insufficient to induce horizons if the scalar potential is too steep to do so Andriot:2023wvg. Nevertheless, this is of course not a proof and further scrutiny may be warranted. Based on these preliminary remarks, we first study spherical ETW branes bounding a solid cylinder of Minkowski space, with the cross section being a ball. Based on the intuition developed, we then proceed by studying a setting where the ETW brane is embedded with flat spatial sections.
@@ 3.1 Spherical ETW Brane
The equations of motion following from the action (9) are [^8] These equations can be obtained from the Israel junction conditions $M_{d}^{d-2}(K_{ab}-h_{ab}K)=S_{ab}$ Israel:1966rt, where $K_{ab}$ is the extrinsic curvature and $S_{ab}$ is the energy momentum tensor of the fields living on the boundary. The extrinsic curvature is defined as $K_{ab}=\nabla_{a}n_{b}$, where $n$ is the outward-pointing unit normal vector of the spacetime boundary.
  Note that our equations differ from much of the brane-cosmology literature (see e.g. Kaloper:1999sm, Binetruy:1999ut, Shiromizu:1999wj, Mizuno:2002wa, Maeda:2000mf) as these papers work locally on the covering space of a $\mathbb{Z}_{2}$ orbifold. By contrast, we treat the brane locally as a boundary of a half-space. Agreement may be achieved through the identification $\rho_{\rm there}=2\rho_{\rm here}$.
$H^{2}=\frac{M_{d}^{-2(d-2)}}{(d-2)^{2}}\rho^{2}-\frac{1}{R^{2}}\,$$\qquad$ $H\equiv\frac{\dot{R}}{R}\,\,\,$, (13)
$M_{d}^{d-3}\left(\ddot{\phi}+(d-2)H\dot{\phi}\right)+\partial_{\phi}V(\phi)=0\,$$\qquad$ $\rho\equiv\frac{1}{2}M_{d}^{d-3}\dot{\phi}^{2}+V(\phi)\,\,$, (14)
and the ‘dot’ stands for $d/d\tau$. Furthermore, it follows from the derivation of (13) that the condition $\rho>0$ must hold. Equation (13) is similar to the Friedmann equation for a closed universe, but with a key difference Binetruy:1999ut, Shiromizu:1999wj: The energy density $\rho$ appears quadratically rather than linearly, implying a substantially different dynamics.
A necessary condition for the existence of brane-brane horizons in the $d$-dimensional spacetime is that the induced, (d1)(d\!-\!1)-dimensional geometry of the boundary has horizons. The condition for these to occur is well-known: It is given by (2) where the scale factor is identified with $R(\tau)$. Note that only the asymptotic form of $R(\tau)$ is of relevance. It follows that if
@@
In the presently discussed situation, the condition above is not sufficient: Indeed, two points $p,p^{\prime}$ on the boundary can be causally connected in the full $d$-dimensional spacetime but not in the (d1)(d\!-\!1)-dimensional geometry of the boundary. A sufficient condition for brane-brane horizons to exist is the following: An observer on the ETW brane should not be able to send a signal to the antipodal point of the spherical brane, not even through the bulk. This corresponds to demanding that
@@
$-\infty<\int^{\infty}\left(\dot{R}-\sqrt{1+\dot{R}^{2}}\right)d\tau\simeq-\frac{1}{2}\int^{\infty}\frac{1}{\dot{R}}d\tau\,$, (19)
@@ 3.2 ETW Branes with flat Slicing
We believe the spherical case discussed above to be the most intuitive and technically simplest. Nevertheless, as just explained, one could suspect that a flat ETW brane embedding leads to a stronger condition. To clarify this, we build on Horowitz:2000ds, but generalize the analysis to arbitrary dimensions. The metric ansatz is
[…]
where u=tzu=t-z, $v=t+z$ and $ds_{d-2}$ is the spatial metric of the ETW brane. The latter is taken to be localized at $z=0$ and the bulk is the region $z>0$. The metric (21) follows from considering an FLRW-type brane embedding and choosing conformal gauge in the directions transverse to $ds_{d-2}^{2}$. The induced metric on the worldvolume of the ETW brane can be written as
@@
with $\rho$ and $p$ being the energy density and pressure on the ETW brane induced by the field $\phi$. We now make an important assumption: We consider $A$ and $B$ to depend on u=tzu=t-z only. One may understand this requirement as choosing the boundary conditions such that no incoming flux of energy hits the ETW brane. Furthermore, this assumption ensures that the bulk geometry is Riemann-flat. It follows that there remains a single nontrivial Einstein equation (25) to be solved. It can be integrated to yield [^10] The integration constant can be absorbed into a redefinition of $u$.
@@
We thus find, in accordance with Shiromizu:1999wj, Maartens:2003tw,
@@
the Friedmann equation of type (13). Because of the flat brane slicing the curvature term $R^{-2}$ is absent. Equation (32) and the scalar field equation (14) determine the evolution of the system. Our assumption that $A$ and $B$ depend only on u=tzu=t-z is related to setting the projected 5d Weyl tensor $E_{ab}$ to zero, see e.g. Shiromizu:1999wj, Maartens:2003tw. We note that the second Israel condition (29) is a consequence of (28) and the the scalar field equation (14).
@@
Importantly, this shows that for $\alpha>2$, the coordinate range of $t$ at the ETW brane is bounded by some constant $u_{0}$. In other words, $\tau\to\infty$ corresponds to $t\to u_{0}$. By contrast, for $1<\alpha<2$, $\tau\to\infty$ corresponds to $t\to\infty$. Furthermore, it is shown in App. A that the functions $e^{A(u)}$ and $e^{B(u)}$ behave as
[…]
@@
$x(t)=\int_{\bar{t}}^{t}e^{(A-B)(\tilde{t}-z(\tilde{t}))}\sqrt{1-z^{\prime}(\tilde{t})}\;d\tilde{t}\,$. (36)
@@
Let us now consider the case $1<\alpha<2$, for which $t$ is unbounded from above. In terms of the variable u(t)=tz(t)u(t)=t-z(t), the integral (36) reads
$x(t)=\int_{\bar{t}}^{u(t)}e^{(A-B)(u)}\frac{\sqrt{(1-z^{\prime})(1+z^{\prime})}}{1-z^{\prime}}du\,.$ (37)
@@
The corresponding coordinate transformation is derived in App. A.2 for the case $1<\alpha<2$. There, it is also shown that the ETW brane embedding is specified by the equation
@@ 3.3 Implications for the Potential
This is known in the case $d=5$, where the corresponding scaling solutions have been studied in the context of brane cosmologies Mizuno:2002wa (see also Maeda:2000mf, Mizuno:2004xj).
@@
Equations (53)–(56) together with the definition (47) are the first consequences of the Refined Horizon Criterion.
@@
As a final comment, let us briefly think about explicit stringy examples for brane-localized scalars: One class of models are those of ‘Verlinde type’ Verlinde:1999fy (mentioned in the Introduction), where a 6d compact space realizes the 4d ETW brane (concretely a $T^{6}/\mathbbm{Z}_{2}$) and a 5d compact space (concretely a fluxed $S^{5}$) realizes the bulk by dimensional reduction. The brane scalar would be the volume of the $T^{6}/\mathbbm{Z}_{2}$. Crucially, if this volume rolls to infinity, higher-dimensional brane-localized operators (in our case a 4d Einstein-Hilbert term) may come to dominate. This situation could arise fairly generally since infinite distances are almost always related to some type decompactification Lee:2019wij. Thus, more work is required to make contact between our simple analysis and string models. Alternatively, if both bulk and brane decompactify or roll to weak coupling together, the above issue is avoided. Thus, we now turn to the case with a bulk scalar field which, as we just argued, may be the more interesting one from the string-theory perspective.
@@ 4 ETW brane coupled to massless Bulk Scalar
As already explained in the Introduction, one may think of $\phi$ as a bulk modulus. The brane generically has less supersymmetry and it is hence natural to expect that it provides a non-trivial potential depending on $\phi$. We do not include a brane-localized kinetic term for $\phi$: While such a term is generically present, its effects are sub-leading Daniel_Master. [^12] On the one hand, brane-parallel spatial gradients decay quickly as the brane expands. On the other hand, the kinetic energy $\dot{\phi}^{2}$ receives contributions from the brane and from the adjacent bulk region, with the latter dominating. This can be traced to a $1/M_{P}$ suppression arising on dimensional grounds Daniel_Master.
@@ 4.1 The Equations of Motion
$\phantom{\Big[}0$ $=$ $\square_{\rm g}\phi\,$, (62)
$0$ $=$ $\left(\left.-M_{d}^{d-2}n^{\mu}\partial_{\mu}\phi-\partial_{\phi}V\right)\right|_{\rm ETW}\,$, (63)
where $\mathcal{K}_{d,\tau\tau}$ is the $\tau\tau$-component of the extrinsic curvature tensor. As before, $\tau$ labels the eigentime of the ETW brane such that $h_{\tau\tau}=-1$. The boundary condition (63) is the straightforward generalization of the familiar Neumann boundary condition to the case with a boundary-localized potential, see e.g. Chamblin:1999ya. In the equations above, $n^{\mu}$ is the (outwards-pointing) unit normal vector of the spacetime boundary and $|_{\rm ETW}$ stands for evaluation at the ETW brane. As before, we focus on positive potentials that approach zero monotonically as $\phi\to\infty$.
We again build on Horowitz:2000ds and consider a metric ansatz of the form (21). As before, we assume the functions $A$ and $B$ as well as the scalar field $\phi$ to depend on u=tzu=t-z only. The single remaining nontrivial Einstein equation now takes the form